OpenStudy (anonymous):

One side of a triangle is 15 inches, and the area of the triangle is 90 sq inches. Find the area of a similar triangle in which the corresponding side is 9 inches. 32.4 sq in 54 sq inch 150 sq inch

OpenStudy (anonymous):

I put 32.4 i just don't know how to get there. The first attempt i put 54 and got it wrong

OpenStudy (driftracer305):

yea..... actually 32.4 is correct...

OpenStudy (anonymous):

How did you get to that answer

OpenStudy (driftracer305):

k watch this.......... what is the formula for area of the first triangle?

OpenStudy (anonymous):

On my lesson thing it says that the formula is area/area= s1/s2

OpenStudy (driftracer305):

k ignore that

OpenStudy (driftracer305):

now look at this.... \[\frac{ 1 }{ 2 } \times base \times 15 = 90\]

OpenStudy (driftracer305):

so, by solving, we get base base as 12

OpenStudy (driftracer305):

now similarly... try finding a corresponding base of other triangle ... \[\frac{ 9 }{ 15} = \frac{ x }{ 12 }\]

OpenStudy (driftracer305):

so u get x as 7.2...........

OpenStudy (anonymous):

alright I got that

OpenStudy (driftracer305):

now plug it in the formula for other triangle,..... \[\frac{ 1 }{ 2 } \times 7.2 \times 9 = 32.4\]

OpenStudy (driftracer305):

u good @Maccrea ?

OpenStudy (anonymous):

Ahh ok I get it now, there are just so many different formulas to remember. Thanks for the help man

OpenStudy (driftracer305):

sure dude.......... anytime :D

OpenStudy (anonymous):

😄